OpenStudy (anonymous):

Write the slope-intercept form of the equation of the line (-5,7) and (4,-2)

OpenStudy (anonymous):

slope = -1 y=-x+2

OpenStudy (anonymous):

use the form y-y1=m(x-x1) find the slope and choose one point and plug in its coordinates